# Thermal Stoner-Wohlfarth Model for Magnetodynamics of Single Domain Nanoparticles: Implementation and Validation

### v1

@@ Abstract
We present the thermal Stoner-Wohlfarth (tSW) model and apply it in the context of Molecular Dynamics simulations. The model is validated against an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles (solid superparamagnet) and a classical dilute (non-interacting) ferrofluid for different combinations of anisotropy strength and magnetic field/moment coupling, at a fixed temperature. We compare analytical and simulation results to quantify the viability of the tSW model in reproducing the equilibrium and dynamic properties of magnetic soft matter systems. We show that if the anisotropy of a particle is more than four-five times higher than the thermal fluctuations, tSW is applicable and efficient. It provides a valuable insight into the interplay between internal magnetization degrees of freedom and Browninan rotation that is often neglected in the fixed point-dipole representation-based magnetic soft matter theoretical investigations.
@@ I Introduction
The idea to control materials using magnetic fields, inspired in the early 20th century by the works of Langevin [1, 2] and through pioneering works on magnetic fluids by Elmore [3] and later by Resler and Rosenzweig [4], has since manifested a branch of science studying magnetic soft matter. The fundamental insight from these pioneering works is that small ferromagnetic particles suspended in a magneto-passive liquid represented a paramagnetic system whose magnetisation can be described with a Langevin function. In other words, one understood that it is possible to engineer and appropriate the magnetic response of complex systems from the collective magnetic response of constituent nanoparticles that themselves have well defined and finely controllable magnetic properties. Under the umbrella term of magnetic soft matter, various magnetoresponsive systems have been explored, including ferrofluids,[5, 6, 7, 8] ferrogels,[9, 10] elastomers,[11, 12, 13, 14, 15] magnetic gels [16, 9, 17] and magnetic filaments (MFs) [18, 19], with a research action focused on generating a macroscopic response by coupling the magnetic response of magnetic nanoparticles, that can be affected via magnetic fields in a contained and precise manner, mechanically or chemically to its environment, such as a polymeric matrix or a viscous environment in general. This class of problem lends itself particularly for classical simulation studies such as Molecular Dynamics (MD). The key that unlocked such simulation studies of composite, complex magnetoresponsive systems, their magnetic properties and response to magnetic fields, is the fixed point dipole representation [20]. In this picture, the dipole moment is depicted as vector that is always coaligned with the anisotropy axis of the nanoparticle, with a fixed magnitude and orientation in the particle reference frame. The fixed point dipole representation of magnetic colloids in general is at the heart of the overwhelming majority of theoretical and classical simulation studies of magnetic fluids and magnetic soft matter [21, 22, 23, 24, 25]. It is a powerful and flexible representation, that albeit undoubtedly practical, incurs important assumptions and approximations. There are no degrees of freedom for the dipole moment to exploit with respect to the anisotropy axis. In other words, magnetic relaxation can only manifest through physical rotation of the nanoparticle body, commonly refereed to as Brownian relaxation. Strictly speaking, the fixed point dipole representation is appropriate for single-domain nanoparticles with infinitely high uniaxial anisotropy [26]. If one is only interested in applied field coupling and far field interactions, more complex magnetic colloids can also be represented as fixed point dipoles, as long as they can be considered as ideally ferro-/ferrimagnetic [27]. Clearly, the fixed point dipole representation fundamentally limits the scope of the studies based on it. In general, depending on the colloids size and/or material, it can neither be taken as given that the dipole moment is coaligned with the anisotropy axis, nor that it has a fixed orientation with respect to it. In fact, the dipole moment has degrees of freedom within the nanoparticle body, which manifest as additional magnetic relaxation mechanisms, most prominent of which is commonly refereed to as Néel relaxation. Generally, the dipole moment states within a nanoparticle correspond to local free energy minima. With Néel relaxation one implies that thermal fluctuations can cause jumps between the available states. In other words, the dipole moment can, driven by thermal fluctuations, flip its orientation with respect to the easy axis. In this context, the anisotropy energy can be interpreted as an energy barrier that needs to be surmounted in order for such a flip to occur. The Brownian and Néel relaxation are two major mechanisms in the description of the magnetic properties of single-domain nanoparticles, and are well captured with a combination of the ideas presented in the seminal works of Stoner and Wohlfarth [28], Néel [29] and Brown [30]. When modeling the dynamics of magnetic soft matter, capturing Brownian and Néel relaxation is fundamentally important [31, 32, 33, 34, 35, 36]. Beyond the academic interest in accurately describing the dynamics of magnetic nanoparticles, understanding magnetic relaxation opens numerous avenues for technological applications [37, 38, 39, 40, 41, 42, 43]. Among others, the interplay between different relaxation mechanisms is key for optimizing magnetic hyperthermia [44, 45, 46, 47].
@@
In this work, we present the thermal Stoner-Wohlfarth (tSW) model as a low computational cost method that allows one to include explicit and accurate magnetodynamics into simulation studies of magnetic soft matter with single domain, uniaxial nanoparticles. In this context, magnetodynamics entails the Brownian and Néel relaxation mechanisms, and the ability of a dipole moment not to be co-aligned with the nanoparticle anisotropy axis in the presence of magnetic and/or dipole fields. We first introduce a formal problem description in Section I, with a detailed discussion on the relevant energy (I.1) and timescales (I.2). In Section II, the model is validated against: (a) an ensemble of immobilized, non-interacting, randomly oriented uniaxial particles (solid superparamagnet) and (b) a classical dilute (non-interacting) ferrofluid. This is equivalent to making the distinction on the basis of whether the nanoparticles exhibit only Néel relaxation mechanisms or both Brownian and Néel relaxation mechanisms, respectively. In both cases, we compare analytical and simulation results to quantify the viability of the thermal SW model in reproducing equilibrium (magnetisation; II.1) and dynamic (susceptibility; II.2) properties of magnetic soft matter systems. A detailed account of the tSW model implementation, simulation details and units is presented in Section I.3. It is important to note that the ideas behind the tSW model have been discussed and used previously [48, 49]. However, a systematic study of its applicability with respect to anisotropy energy/temperature and magnetic field strength has remained unclear up to now. Moreover, the viability and applicability of the model in the context of MD simulations of magnetic soft matter has not been explored so far.
@@ I.1 Formal problem description
Above the critical field, only one global minimum remains. The particle dipole moment follows the position of the local energy minimum and instantaneously changes its direction depending on strength of $h$ in relation to $h_{cr}$. Transitions between the energy minima are not possible below $h_{cr}$ (at $T=0$). A visualisation of the energy landscape defined with Eq. 5 ($\phi=\zeta$), along with its dependence on $\theta$,$\psi$ and $h$, is provided in Fig.2. The simplest thermal formulation of the total magnetic energy would include the Néel relaxation mechanism, understood as the internal relaxation achieved through thermal fluctuation induced transitions of the dipole moment between the local energy minima in the total magnetic energy of a particle. This intuitive notion of the Néel mechanism inheres several important approximations, discussed at length below. Brownian relaxation instead refers to rotational diffusion of the particle body.
Figure 2: Visualisation of the energy (density) landscape in the Stoner-Wohlfarth model, as a function of $\theta$. Lines with a different colour saturation correspond to energy densities for a different $\psi$. The more saturated a colour appears, the more aligned the anisotropy axis is with the applied field (smaller $\psi$). Different colours correspond to different applied field strengths, where green energy density profiles correspond to a higher applied field strength then blue energy density profiles.)
@@ I.2 Relevant time scales
Here, we discuss the timescales of the relaxation mechanisms characteristic for ensembles of magnetic nanoparticles, starting with the ones associated with internal degrees of freedom. In the absence of thermal fluctuations and an applied field, the main time scale that characterizes the magnetization relaxation in immobilized ensembles, is the damping time of the Larmor precession. From the standard Landau-Lifshitz-Gilbert (LLG) equation, it can be estimated as [50]:
@@
Formally, the Néel time $\tau_{N}$ is the inverse of the smallest non-vanishing eigenvalue of the Fokker-Planck equation, that Brown formulated on the basis of the stochastic Landau-Lifshitz-Gilbert equation. Eq. 9 is only valid for uniaxial particles in zero field. However, in the literature one can find multiple approximations for $\tau_{N}$ in static fields of different orientations, as well as for other anisotropy types [26, 53]. Brown in his work also demonstrated that the zero-field rate of over-barrier jumps is not affected by the Larmor precession and that corresponding term in the Fokker-Planck equation can be safely omitted as one is looking for $\tau_{N}$. The same is true, if the field is applied parallel to the easy axis ($\psi=0$). This is a general property of the Fokker-Plank-Brown equation for problems that possess axial symmetry. However, for all other cases ($\psi\neq 0$ or different anisotropy types), precession will affect magnetic relaxation and should in principle be explicitly taken into consideration [54, 55]. Omitting precession in the general case limits the results to the field frequencies below the ferromagentic resonance range.
@@
where $J=0.1\rho Vd^{2}$ is the moment of inertia, $\rho$ is the particle material density (here we assume, that all the mass is concentrated in the magnetic core), $\Gamma_{R}=\pi\eta d^{3}_{H}$ is the rotational friction coefficient, $\eta$ is the fluid viscosity. The rotational Brownian time is:
@@ I.3 Implementation details
Here we outline our tSW algorithm, specifically, how we simulate the internal dynamics and the Néel relaxation mechanism in magnetic colloids. For more details on the simulation approach, we refer the reader to Section IV. The algorithm requires $\mu$, $h$ and $\sigma$ as input parameters. The algorithm can be logically separated into three steps:
@@
Firstly, $h_{cr}$ (Eq. 6) is calculated taking the current angle between the vector $h$ and the particle anisotropy axis at its position, $\phi$. For a given $\phi$, the algorithm finds $\theta$ which minimises the total magnetic energy (Eq. 5) $\theta^{\prime}_{min}$, which is closest to the previous dipole moment state. This is ensured by initialising the state of the energy minimiser with the previous particle state. If the field acting on the particle is less than $h_{cr}$, the algorithm proceeds to find a $\theta$ that maximises the total magnetic energy from both sides of $\theta^{\prime}_{min}$, denoted with $\theta^{\prime}_{max}$ and $\theta^{\prime\prime}_{max}$. With this information, the algorithm calculates the energy barriers found on both sides of $\theta^{\prime}_{min}$,
@@
and from it, the transition probability $p=1-e^{\delta t/\tau_{N}}$, where $\delta t$ is the integration step time [49]. It is suggested in the literature, for example in Chuev and Hesse [48], that one should use the $\Delta E^{eff}=(\Delta E^{\prime}+\Delta E^{\prime\prime})/2$ barrier, to better approximate the three-dimensional character of the energy barrier and get a better estimate of the transitions probabilities. In our testing, however, we did not find measurable differences when using $\Delta E^{eff}$, and therefore choose the simpler approach. We make a trial move by casting a random number and comparing against the transition probability, akin to a Metropolis step. If the trial move is successful, the algorithm finds a new $\theta$ that minimises Eq. 5, denoted by $\theta^{\prime\prime}_{min}$, and sets the dipole moment to point in accordance with new minimum. Alternatively, the dipole moment is set to point in accordance with $\theta^{\prime}_{min}$.
@@ II Results and Discussion
Looking at Fig. 3, one can get a sense of the considerations and constraints that one comes across in simulations studies of magnetic soft matter, where often, the objective is to scrutinise collective properties at long timescales. In this Figure, we plot the dependence of Neél ($\tau_{N}$), Brownian ($\tau_{B}$), internal diffusion ($\tau_{D}$), period of the precession ($2\pi/\omega_{L}$), precession damping ($\tau_{0}$) and inertial decay ($\tau_{j}$) times as functions of magnetite particle diameter $d$. The range of the anisotropy constant $\sigma$ between $5-20$ covers the most commonly studied nanoparticle sizes and materials. If one was to simulate the magnetodynamics of the nanoparticles explicitly, accounting for all their characteristic magnetic relaxation timescales by, for example, using the aforementioned LLG equation [57, 58, 59], it becomes apparent that $\tau_{J}$ is not small enough to be safely ignored. This means, that one must resolve timescales less than picoseconds to be able to include full magnetodynamics in simulation. This is prohibitively low for any simulations studies concerned with bulk properties. So, from practical considerations, one would like to simulate timescales where high frequency processes ($\leq 10^{-8}$) can be safely ignored. However, a competition remains between $\tau_{N}$ and $\tau_{D}$ on the internal relaxation, and $\tau_{B}$ on the external relaxation side. Specifically, within the $205<\sigma<$ 20 range, there are regions where $\tau_{N}$ and $\tau_{D}$ are comparable within simulation time, to $\tau_{N}$ becoming prohibitively large, if one was to consider relaxation associated with $\tau_{D}$. Understanding the optimal approach to be able to include relevant magnetodynamic processes in this range, is precisely the question we address below.
Figure 4: The absolute magnetisation error $\Delta m$ (in percent) between the tSW model (theory–solid line; simulation–symbols) and continuous model descriptions for a solid superparamagnet (subfigure (a)) or a non-interacting ferrofluid (subfigure (b)), as a function of the Langevin parameter $\xi$ and the anisotropy parameter $\sigma$. The colour mapping for $\sigma$ values explored is provided with a colour-bar.

### v2

@@ Abstract
We present the thermal Stoner-Wohlfarth (tSW) model and apply it in the context of Molecular Dynamics simulations. The model is validated against an ensemble of immobilized, randomly oriented uniaxial particles (solid superparamagnet) and a classical dilute ferrofluid for different combinations of anisotropy strength and magnetic field/moment coupling, at a fixed temperature. We compare analytical and simulation results to quantify the viability of the tSW model in reproducing the equilibrium (with and without dipole-dipole interactions) and dynamic (without dipole-dipole interactions) properties of magnetic soft matter systems. We show that if the anisotropy of a particle is more than five times higher than the thermal fluctuations, tSW is applicable and efficient. This approach allows one to consider the interplay between Néel and Brownian relaxation, often neglected in the fixed point-dipole representation-based magnetic soft matter theoretical investigations.
@@ I Introduction
The idea to control materials using magnetic fields, inspired in the early 20th century by the works of Langevin [1, 2] and through pioneering works on magnetic fluids by Elmore [3] and later by Resler and Rosenzweig [4], has since manifested a branch of science studying magnetic soft matter. The fundamental insight from these pioneering works is that small ferromagnetic particles suspended in a magneto-passive liquid represented a paramagnetic system whose magnetization can be described with a Langevin function. In other words, one understood that it is possible to engineer and appropriate the magnetic response of complex systems from the collective magnetic response of constituent nanoparticles that themselves have well defined and finely controllable magnetic properties. Under the umbrella term of magnetic soft matter, various magnetoresponsive systems have been explored, including ferrofluids,[5, 6, 7, 8] ferrogels,[9, 10] elastomers,[11, 12, 13, 14, 15] magnetic gels [16, 9, 17] and magnetic filaments (MFs) [18, 19], with a research action focused on generating a macroscopic response by coupling the magnetic response of magnetic nanoparticles, that can be affected via magnetic fields in a contained and precise manner, mechanically or chemically to its environment, such as a polymeric matrix or a viscous environment in general. This class of problem lends itself particularly for classical simulation studies such as Molecular Dynamics (MD). The key that unlocked such simulation studies of composite, complex magnetoresponsive systems, their magnetic properties and response to magnetic fields, is the fixed point dipole representation [20]. In this picture, the dipole moment is depicted as vector that is always coaligned with the anisotropy axis of the nanoparticle, with a fixed magnitude and orientation in the particle reference frame. The fixed point dipole representation of magnetic colloids in general is at the heart of the overwhelming majority of theoretical and classical simulation studies of magnetic fluids and magnetic soft matter [21, 22, 23, 24, 25]. It is a powerful and flexible representation, that albeit undoubtedly practical, incurs important assumptions and approximations. There are no degrees of freedom for the dipole moment to exploit with respect to the anisotropy axis. In other words, magnetic relaxation can only manifest through physical rotation of the nanoparticle body, commonly refereed to as Brownian relaxation. Strictly speaking, the fixed point dipole representation is appropriate for single-domain nanoparticles with infinitely high uniaxial anisotropy [26]. If one is only interested in applied field coupling and far field interactions, more complex magnetic colloids can also be represented as fixed point dipoles, as long as they can be considered as ideally ferro-/ferrimagnetic [27]. Clearly, the fixed point dipole representation fundamentally limits the scope of the studies based on it. In general, depending on the colloids size and/or material, it can neither be taken as given that the dipole moment is coaligned with the anisotropy axis, nor that it has a fixed orientation with respect to it. In fact, the dipole moment has degrees of freedom within the nanoparticle body, which manifest as additional magnetic relaxation mechanisms, most prominent of which is commonly refereed to as Néel relaxation. Generally, the dipole moment states within a nanoparticle correspond to local free energy minima. With Néel relaxation one implies that thermal fluctuations can cause jumps between the available states. In other words, the dipole moment can, driven by thermal fluctuations, flip its orientation with respect to the easy axis. In this context, the anisotropy energy can be interpreted as an energy barrier that needs to be surmounted in order for such a flip to occur. The Brownian and Néel relaxation are two major mechanisms in the description of the magnetic properties of single-domain nanoparticles, and are well captured with a combination of the ideas presented in the seminal works of Stoner and Wohlfarth [28], Néel [29] and Brown [30]. When modeling the dynamics of magnetic soft matter, capturing Brownian and Néel relaxation is fundamentally important [31, 32, 33, 34, 35, 36]. Beyond the academic interest in accurately describing the dynamics of magnetic nanoparticles, understanding magnetic relaxation opens numerous avenues for technological applications [37, 38, 39, 40, 41, 42, 43]. Among others, the interplay between different relaxation mechanisms is key for optimizing magnetic hyperthermia [44, 45, 46, 47].
In this work, we present the thermal Stoner-Wohlfarth (tSW) model as a low computational cost method that allows one to include explicit and accurate magnetodynamics into simulation studies of magnetic soft matter with single domain, uniaxial nanoparticles. In this context, magnetodynamics entails: (i) Brownian relaxation; (ii) Néel relaxation mechanisms, and (iii) the ability of a dipole moment not to be co-aligned with the nanoparticle anisotropy axis in the presence of magnetic and/or dipole fields. Talking about low computational cost methods to include explicit magnetodynamics, one must mention the highly efficient and accurate approaches to numerical solve of the Fokker-Planck equation [48, 49], for the so-called “Egg model” of superparamagnetic particle dynamics initially developed by Shliomis and Stepanov [50]. While solving the Fokker-Planck equation is a very powerful and fast approach to solve magnetodynamics, it is restricted to relatively simple non-interacting systems, and in general, is less flexible than stochastic simulations. It is also worth highlighting the diffusion jump model [51, 52], as an example of a similar, low computational cost method to simulate magnetodynamics. This model however, does not capture the (iii) aspect of magnetodynamics present in the tSW model, because in it, the magnetic moment is necessarily aligned with the crystallographic axis of the colloid.
@@
The paper is structured as follows: We first introduce a formal problem description in Section I, with a detailed discussion on the relevant energy (I.1) and timescales (I.2). In Section II, the model is validated against: (a) an ensemble of immobilized, randomly oriented uniaxial particles (solid superparamagnet) and (b) a classical dilute ferrofluid. This is equivalent to making the distinction on the basis of whether the nanoparticles exhibit only Néel relaxation mechanisms or both Brownian and Néel relaxation mechanisms, respectively. In both cases, we compare analytical and simulation results to quantify the viability of the thermal SW model in reproducing equilibrium (magnetization for a non-interacting and interacting system; II.1) and dynamic (susceptibility for a non-interacting system; II.2) properties of magnetic soft matter systems. Note that throughout the manuscript, unless explicitly stated otherwise, we are referring to non-interacting systems. A detailed account of the tSW model implementation, simulation details and units is presented in Section I.3. The theoretical framework behind the tSW model has been introduced in Chuev and Hesse [53], where it is called the generalized SW model. The generalized SW model has been implemented and used in several studies under the name of kinetic Monte Carlo (kMC) approach. [54, 55, 56, 57, 58] However, the kMC terminology has been used inconsistently in the literature. It is important to clarify that, broadly speaking, the kMC approach implies that the magnetic degrees of freedom are governed by sampling the magnetic energy in the SW model via a MC algorithm. The transition probabilities are calculated in accordance with the generalized SW model. The kMC approach is, however, not limited to a two state approximation. In other words, in the kMC approach, the dipole moment does not need to point in accordance with the minima of the magnetic energy. Therefore, it should be considered as distinct from both the generalized SW and the tSW model. The tSW approach can be seen as a subset of kMC, in the sense that it strictly adheres to the two state approximation. However, the tSW model can also be seen as a generalization of the kMC approach, in the sense that couples it with Langevin Dynamics. Regardless, in this contribution, we present the first systematic study of the applicability of the tSW model with respect to anisotropy energy/temperature and magnetic field strength, where the viability and applicability of the model in the context of MD simulations of magnetic soft matter is qualified.
@@ I.1 Formal problem description
Above the critical field, only one global minimum remains. The particle dipole moment follows the position of the local energy minimum and instantaneously changes its direction depending on the strength of $h$ in relation to $h_{cr}$. Transitions between the energy minima are not possible below $h_{cr}$ (at $T=0$). A visualization of the energy landscape defined with Eq. 5 ($\phi=\zeta$), along with its dependence on $\theta$,$\psi$ and $h$, is provided in Fig.2. The simplest thermal formulation of the total magnetic energy would include the Néel relaxation mechanism, understood as the internal relaxation achieved through thermal fluctuation induced transitions of the dipole moment between the local energy minima in the total magnetic energy of a particle. This intuitive notion of the Néel mechanism involves several important approximations, discussed at length below. Brownian relaxation instead refers to the rotational diffusion of the particle body.
Figure 2: Visualization of the energy (density) landscape in the Stoner-Wohlfarth model, as a function of $\theta$. Lines with a different color saturation correspond to energy densities for a different $\psi$. The more saturated a color appears, the more aligned the anisotropy axis is with the applied field (smaller $\psi$). Different colors correspond to different applied field strengths, where green energy density profiles correspond to a higher applied field strength then blue energy density profiles.)
@@ I.2 Relevant time scales
Here, we discuss the timescales of the relaxation mechanisms characteristic for ensembles of magnetic nanoparticles, starting with the ones associated with internal degrees of freedom. In the absence of thermal fluctuations and an applied field, the main timescale that characterizes the magnetization relaxation in immobilized ensembles is the damping time of the Larmor precession. From the standard Landau-Lifshitz-Gilbert (LLG) equation, it can be estimated as [59]:
@@
Formally, the Néel time $\tau_{N}$ is the inverse of the smallest non-vanishing eigenvalue of the Fokker-Planck equation, that Brown formulated on the basis of the stochastic LLG equation. Eq. 9 is only valid for uniaxial particles in zero field. However, in the literature one can find multiple approximations for $\tau_{N}$ in static fields of different orientations, as well as for other anisotropy types [26, 62]. Brown, in his work, also demonstrated that the zero-field rate of over-barrier jumps is not affected by the Larmor precession and that the corresponding term in the Fokker-Planck equation can be safely omitted when one is looking for $\tau_{N}$. The same is true if the field is applied parallel to the easy axis ($\psi=0$). This is a general property of the Fokker-Plank-Brown equation for problems that possess axial symmetry. However, for all other cases ($\psi\neq 0$ or different anisotropy types), precession will affect magnetic relaxation and should in principle be explicitly taken into consideration [63, 64]. Omitting precession in the general case limits the results to the field frequencies below the ferromagnetic resonance range.
@@
where $J=0.1\rho Vd^{2}$ is the moment of inertia, $\rho$ is the particle material density (here we assume that all the mass is concentrated in the magnetic core), $\Gamma_{R}=\pi\eta d^{3}_{H}$ is the rotational friction coefficient, $\eta$ is the fluid viscosity. The rotational Brownian time is:
@@ I.3 Implementation details
Here we outline our tSW algorithm, specifically, how we simulate the internal dynamics and the Néel relaxation mechanism in magnetic colloids. For more details on the simulation approach, we refer the reader to Section IV. The algorithm requires $\mu$, $h$, and $\sigma$ as input parameters. The algorithm can be logically separated into three steps:
@@
Firstly, $h_{cr}$ (Eq. 6) is calculated taking the current angle between the vector $h$ and the particle anisotropy axis at its position, $\phi$. For a given $\phi$, the algorithm finds $\theta$ which minimizes the total magnetic energy (Eq. 5) $\theta^{\prime}_{min}$, which is closest to the previous dipole moment state. This is ensured by initializing the state of the energy minimizer with the previous particle state. If the field acting on the particle is less than $h_{cr}$, the algorithm proceeds to find a $\theta$ that maximizes the total magnetic energy from both sides of $\theta^{\prime}_{min}$, denoted with $\theta^{\prime}_{max}$ and $\theta^{\prime\prime}_{max}$. With this information, the algorithm calculates the energy barriers found on both sides of $\theta^{\prime}_{min}$,
@@
and from it, the transition probability $p=1-e^{\delta t/\tau_{N}}$, where $\delta t$ is the integration step time [54]. It is suggested in the literature, for example, in Chuev and Hesse [53], that one should estimate the transition probability as an average, considering both $\Delta E^{\prime}$ and $\Delta E^{\prime\prime}$. While the general mean barrier approach is a topic for discussion [66], in our testing, we did not find a measurable difference between the two approaches and used the simpler one that is compatible with the analytical formulation of tSW. We make a trial move by casting a random number and comparing it against the transition probability, akin to a Metropolis step. If the trial move is successful, the algorithm finds a new $\theta$ that minimizes Eq. 5, denoted by $\theta^{\prime\prime}_{min}$, and sets the dipole moment to point in accordance with the new minimum. Alternatively, the dipole moment is set to point in accordance with $\theta^{\prime}_{min}$.
@@ II Results and Discussion
Looking at Fig. 3, one can get a sense of the considerations and constraints encountered in simulation studies of magnetic soft matter, where the objective is often to scrutinize collective properties at long timescales. We plot the dependence of Neél ($\tau_{N}$), Brownian ($\tau_{B}$), internal diffusion ($\tau_{D}$), period of the precession ($2\pi/\omega_{L}$), precession damping ($\tau_{0}$) and inertial decay ($\tau_{j}$) times as functions of magnetite particle diameter $d$. The range of the anisotropy constant $\sigma$ between $5-20$ covers the most commonly studied nanoparticle sizes and materials. If one was to simulate the magnetodynamics of the nanoparticles explicitly, accounting for all their characteristic magnetic relaxation timescales by, for example, using the aforementioned LLG equation [67, 68, 69], it becomes apparent that $\tau_{J}$ is not small enough to be safely ignored. This means that one must resolve timescales less than picoseconds to be able to include full magnetodynamics in simulation. This is prohibitively low for any simulation study concerned with bulk properties. So, from practical considerations, one would like to simulate timescales where high frequency processes ($\leq 10^{-8}$) can be safely ignored. However, a competition remains between $\tau_{N}$ and $\tau_{D}$ on the internal relaxation, and $\tau_{B}$ on the external relaxation side. Specifically, within the $205<\sigma<$ 20 range, there are regions where $\tau_{N}$ and $\tau_{D}$ are comparable within simulation time, to $\tau_{N}$ becoming prohibitively large, if one was to consider relaxation associated with $\tau_{D}$. Understanding the optimal approach to be able to include relevant magnetodynamic processes in this range, is precisely the question we address below.
Figure 4: The absolute magnetization error $\Delta m$ (in percent), as a function of $\xi$, for a non-interacting (a) solid superparamagnet, where the analytical tSW results (solid lines, given by Eq. 15) and simulations (symbols) are compared to the continuum-level description (given by Eq. 45); (b) ferrofluid, where the analytical tSW results (solid lines, given by Eq. 17) and simulations (symbols) are compared to the continuum-level description (given by Eq. 18). The color bar indicates the anisotropy parameter $\sigma$ values explored.